OpenStudy (anonymous):

2x-6=4x+2 Someone help please will give Medal

OpenStudy (precal):

move the x's to one side by taking the opposite sign and move the number to the other side

OpenStudy (precal):

you need to isolate your x variable to find its value

OpenStudy (anonymous):

Ok thank you!

OpenStudy (precal):

do you want me to check your work

OpenStudy (anonymous):

So -6=2x+2 (-6-2)=-8 -8=2x divide the 2 to -8 and x=-4

OpenStudy (precal):

2x-6=4x+2 -2x -2x -6=2x+2 -2 -2 -8=2x -4=x yes

OpenStudy (anonymous):

Alright thank you for your help!

OpenStudy (precal):

anytime